A body dropped from a tall building covers,two floors in the first second.If all the floors are of equal height,the number of floors it covers in the next one second is

Grade:11

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Chakrala,

 

Let height of 1 floor is h,

u = 0 , so,  v  = u + at     (a = g =10 m/s2 )

                 v = 10 * 1 = 10 m/s

  Now,    v2 - u2 = 2as

 so,     100  = 2 * 10 * s

    or,   s = 5 m

which implies height of each floor is  = 5 / 2 =  2.5 m

For 2 sec. , distance covered is  s= ut + 1/2 at2 = 0 + 10 * 4/2 = 20m

So, we have Distance covered in 2nd second = 20 - 5 = 15 m

Henc e, No. of floor = 15/2.5 = 6

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay